OpenStudy (undeaddragon26):

Which equivalence factor should you use to convert from 6.26 moles of O2 to molecules of O2? Answer (1 molecule/6.02 x 1023 mol O2) (6.02 x 1023 molecules O2/1 mol O2) (6.02 x 1023 molecules O2/6.26 mol O2) (6.26 mol O2/6.02 x 1023 molecules O2)

OpenStudy (abmon98):

none of them are right check your question. 1 mole contains 6.02*10^23 molecules (Avogadro s constant number) 6.26 moles contains x molecules x=6.26*6.02*10^23/1

OpenStudy (somy):

you can use (6.02 x 1023 molecules O2/1 mol O2)

OpenStudy (somy):

\[\frac{ 6.02 \times 10^23 ~ molecules~of~O_2 }{ 1 ~mole~ of~ O_2 } = \frac{ X~molecules~of~O_2 }{ 6.26~moles~of~O_2 }\]

OpenStudy (somy):

btw @Abmon98 read the question again lol they are saying 'equivalence factor'

OpenStudy (somy):

its more like math tho

OpenStudy (abmon98):

sorry i misunderstood the question @somy and keep it up ;)

OpenStudy (somy):

yeah the wording is not good so its not ur fault <.< i got it from the way answers look like